OpenStudy (anonymous):

Sketch the indicated and integrate f(x,y) over D using polar coordinates. f(x,y)=2xy x>=0 , y>=0, x^2+y^2<=49 ;

OpenStudy (anonymous):

I got 7, where did i go wrong???

OpenStudy (anonymous):

Dabears :D After I grab a shower I'll help you :P Give me 10 :P

OpenStudy (anonymous):

yay! thanks malevolence :)

OpenStudy (anonymous):

x^2 +y^2 <=49 will give you in polar coordinates: r<=7 right?

OpenStudy (anonymous):

\[x=r*\cos \alpha\] \[y=r*\sin \alpha\] integrate

OpenStudy (anonymous):

yes thats why I was thinking 7ish?

OpenStudy (anonymous):

\[=\int\limits_{0}^{\pi/2}\int\limits_{0}^{7}2*r ^{2}*\sin \alpha *\cos \alpha drd \alpha\] seems like this should be right for x>=0 & y>=0

OpenStudy (anonymous):

ohhh i think i see what i did

OpenStudy (anonymous):

which gives me: \[343/3 \int\limits_{0}^{\pi/2} \sin2\alpha d \alpha=343/3 (-\cos2\alpha)/2=343/3\]

OpenStudy (anonymous):

how we doing? does it match the answer?

OpenStudy (anonymous):

let me plug it in real quick

OpenStudy (anonymous):

said it was wrong :(

OpenStudy (anonymous):

shoot.... let me think

OpenStudy (anonymous):

maybe for the polar coordinates ? f(x,y) = 2xy = 2(rcos (theta)(r sin theta) = r^2 sin 2 theta

OpenStudy (anonymous):

I did the same thing...exactly

OpenStudy (anonymous):

how did you get the cos for the integral above? I used same intervals but had (r^2sin2theta) no cos()

OpenStudy (anonymous):

(r^2sin2theta)r dr dtheta -- i mean

OpenStudy (anonymous):

you have the same... actually x=rcos 7 y=rsin so, F=2r^2 sin * cos = r^2 sin2 theta

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin2\beta d \beta=-1/2 \cos2\beta\] right?

OpenStudy (anonymous):

oh yeah, whooops :)

OpenStudy (anonymous):

i do not see any mistakes... I keep getting 343/3 as an answer... I have a special software! - let me check!

OpenStudy (anonymous):

okay thanks !

OpenStudy (anonymous):

Well you have: \[\int\limits_{0}^{\frac{\pi}{2}} \int\limits_{0}^{7}2r^2\sin(\theta)\cos(\theta)r dr d \theta\] So that gives r^3. Which makes the integrand r^3sin(2theta)

OpenStudy (anonymous):

Integrating r you get (1/4)r^4sin(2theta) Or: (2401/4)sin(2theta) Then integrating theta you get: (2401/4)(-1/2)cos(2theta) Evaluating: (-2401/8)(-1-1)=2401/4?

OpenStudy (anonymous):

Which is what wolfram gives me.

OpenStudy (anonymous):

if this wasn't the internet, I'd shake your hand! thank you it was right

OpenStudy (anonymous):

I forgot that pesky r on the outside!!

OpenStudy (anonymous):

I even wrote it above haha

OpenStudy (anonymous):

:D I'm glad. Mathematica also confirms this :P (wolfram but with straight input code)

OpenStudy (anonymous):

perfect thanks again! and thx inik for helping me too!

OpenStudy (anonymous):

I missed r^4.... ! shame :(